A 20 Kg block rest on horizontal frictionless surface. a constant force of 50 N is applied for 3 seconds. What is the impulse acting on the block? What is the final speed of the block?
1
Expert's answer
2020-12-02T07:15:32-0500

"p=mv=mat=Ft=50\cdot3=150 ~\frac{kg\cdot m}{s},"

"v=at=\frac Fm\cdot t=\frac{50}{20}\cdot 3=7.5 ~\frac ms."